OpenStudy (anonymous):

2. Solve for x: 2x−11=7x+19 A. x = −6 B. x = 8/9 C. x = 10/3 D. x = −8/5

OpenStudy (anonymous):

2x - 11 = 7x -19 -19 - 11 = 7x - 2x -30 = 5x x = -30/5 = -6

OpenStudy (anonymous):

x = -6

OpenStudy (anonymous):

2x−11=7x+19 2x-7x=19+11 -5x=30 x=30\-5 x=-6